OpenStudy (anonymous):

If a baseball team won 98 games and lost 64 games, what percent of games did they win? (Round to two decimal places as needed)

OpenStudy (danjs):

Wins / (Wins+Loss) or wins/total games

OpenStudy (anonymous):

so 162/ 98?

OpenStudy (anonymous):

sorry 98/162

OpenStudy (anonymous):

which equals 60.50%?

OpenStudy (danjs):

they won 98 out of (98+64) games 98/(98+64) or 98/162

OpenStudy (danjs):

that is a decimal value out of 1 percent is out of 100, so multiply by 100/100 60.5/100, , or you can say 60.5% per'cent' --- per one hundred

OpenStudy (anonymous):

see that's what I thought the answer was being as it said, "Round to two decimal places as needed" but then I typed in 60.50% and it said that it was wrong, that the correct answer was 60.49%. How could that be?

OpenStudy (danjs):

yeah that 98/162 = 0.60493 the percent is 100 times that, 60.493% round to 2 decimals ----60.49%

OpenStudy (danjs):

sorry i just used bantaray value, didnt calculate it myself

OpenStudy (anonymous):

oh got it! I was only rounding one place. thanks

OpenStudy (danjs):

you can always just figure 98/162 = x/100 x is the percent value 98 out of 162 is the same as 'what number' over 100